OpenStudy (anonymous):

Balls are randomly removed from a bag without replacement. If the probability that the first 5 balls drawn from the bag are all green is 1/2, what is the least possible number of balls in the bag at the start? the answer would be 10, right?

OpenStudy (anonymous):

I believe 10 is correct

OpenStudy (anonymous):

!!! thx... can you help me with something else?

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

What ordered triple of integers (a,b,c), with c>b>a>0, satisfies (root a + root b + root c)^2=10+ root 24 +root 40 + root 60 ?

OpenStudy (anonymous):

which, i know the last three roots can be simplified, and i finally managed to remember (i think) how to simplify that whole left side, so now i have a+b+c+2rootab+2rootbc+2rootac+20+2root6+2root10+2root15

OpenStudy (anonymous):

btw im having trouble typing, so i type stuff out the long way like roots... sorry about that

OpenStudy (anonymous):

and i was thinking... ab, bc, and ac could be in some order equal to 6, 10, and 15, couldn't they? so wouldnt i make a system of equations for that?

OpenStudy (anonymous):

sorry ... i got pulled away from the computer...this question is actually pretty easy

OpenStudy (anonymous):

\[(\sqrt{a} + \sqrt{b} + \sqrt{c})^2=10+\sqrt{24} +\sqrt{40} + \sqrt{60}\] \[a+b+c+2(\sqrt{ab}+\sqrt{ac}+\sqrt{bc})=10+2\sqrt{6} +2\sqrt{10} + 2\sqrt{15}\] \[a+b+c+2(\sqrt{ab}+\sqrt{ac}+\sqrt{bc})=10+2(\sqrt{6} +\sqrt{10} + \sqrt{15})\] so we can have \(a+b+c=10\) and \(ab=6, ac=10\) and \(bc=15\) a solution is \(a=2, ~b=3\) and \(c=5\)

OpenStudy (anonymous):

sorry, i left too... so basically there's no system involved, really, you just... oh. ok. i see now. thanks!!!

OpenStudy (anonymous):

umm, how did you guys end up with 10?

OpenStudy (anonymous):

um... was i anywhere near the right track previously?

OpenStudy (anonymous):

oh, sorry, kevin person, that was a different problem... are you talking about the balls one? scrath the "different problem"... ok...

OpenStudy (anonymous):

so i think it's because the probability of there being all green balls in there five drawn is one half, you would double the amount of balls at hand to get the total amount of balls.

OpenStudy (anonymous):

i think.

OpenStudy (anonymous):

from the 10 balls in the bag originally, how many of them are green?

OpenStudy (anonymous):

five... the problem said that, i thought...

OpenStudy (anonymous):

waaaiiiiit....

OpenStudy (anonymous):

then, probability of first five balls being all green = 5/10*4/10*3/10*2/10*1/10 which clearly is not 1/2

OpenStudy (anonymous):

ok. so then what'd you do? i am perfectly helpless when it comes to probability, so it looks like i need instruction

OpenStudy (anonymous):

well I just know the answer is not 10, I'm still thinking about it, gimme a sec

OpenStudy (anonymous):

I'm sorry the above should be 5/10*4/9*3/8*2/6*1/5

OpenStudy (anonymous):

wow it's 10 balls with 9 of them being green. Nice coincidence!! Then the required probability would be\[P=\frac{ 9 }{ 10 }\times \frac{ 8 }{ 9 }\times \frac{ 7 }{ 8 }\times \frac{ 6 }{ 7 }\times \frac{ 5 }{ 6 }\] which simplifies nicely to 5/10 which is 1/2

OpenStudy (anonymous):

wait, how did you do that? first off, how'd you figure 9 out of 10? or did you just do the above work backwards? that cant be what you did... what did you do?

OpenStudy (anonymous):

I really don't know! The pattern of cancelling the numerator of one term with the denominator of the other came to my mind and I chose the first term such that the denominator of the first term would be twice the numerator of the last.

OpenStudy (anonymous):

there are five balls so there should be five terms

OpenStudy (anonymous):

well... ok, that makes sense... but how did you get those terms?! HOW DID YOU COME UP WITH 9 OUT OF 10?!

OpenStudy (anonymous):

so 10 and 5 should be the denominator of the first term and numerator of the last term, right? Since I need numerator of every term to cancel the denominator of the next term I had to choose 9, since removing one ball would mean the denominator of the second term is 9

OpenStudy (anonymous):

ohhh... ok, i get that. ok, thanks! makes better sense, now, i think... so the answer IS 10? or isnt it?

OpenStudy (anonymous):

yeah, it seems so

OpenStudy (anonymous):

ok, cool. so i was right before, just did it wrong. lucky.